# The Reality of AI and Biorisk

### v2

@@ 3.2.2 Relevant research
What does this mean for predicting unknown toxins or proteins that underpin harmful biological artifacts? Designing novel diseases or bioweapons is considerably difficult (European Commission, 2005; National Academies Press, 2004). Skilled and technical expertise is required to utilize BTs effectively, such as cell culturing skills or experience working with genomes, in addition to hard-to-obtain or expensive material as well as organizational and staff resources (National Academies Press, 2004; Ouagrham-Gormley, 2014) (See Table 2). This suggests that these risks are primarily limited to state actors or sophisticated users with advanced knowledge in biology and machine learning, reducing both scale of risk and the “attack surface” that needs to be monitored (Rose et al., 2024; Ouagrham-Gormley, 2014).
@@ 4 Other potential threat models
Other, analyses highlight that LLMs may considerably or significantly uplift malicious users’ ability to develop harmful biological artifacts through improving laboratory experimentation and troubleshooting, and reducing barriers to biological work, for example by enabling technical coding for computational biology to be performed with non-technical, natural language (Rose et al., 2024). There are few empirical studies of this threat model, though one example is reported in OpenAI’s o1 system card which assesses model capabilities across a range of biological tasks, such as troubleshooting wet lab protocols and automating wet lab work. From these evaluations, OpenAI reports that o1-preview and o1-mini pose “limited risk,” because the models do not provide biology experts with new knowledge, and "do not enable non-experts to create biological threats, because creating such a threat requires hands-on laboratory skills that the models cannot replace” (OpenAI, 2024c). With regards to AI tools assisting lab-based work, it is notable that existing risk frameworks focus on mitigating physical synthesis of threats, for example focusing on containment and control procedures that secure the handling and production of biological artifacts, rather than digital capabilities that support knowledge gathering, task automation, or planning (US Department of Health and Human Sciences, 2015).

### v3

@@ 3.2.2 Relevant research
What does this mean for predicting unknown toxins or proteins that underpin harmful biological artifacts? Designing novel diseases or bioweapons is considerably difficult (European Commission, 2005; National Academies Press, 2004). Skilled and technical expertise is required to utilize BTs effectively, such as cell culturing skills or experience working with genomes, in addition to hard-to-obtain or expensive material as well as organizational and staff resources (National Academies Press, 2004; Ouagrham-Gormley, 2014) (See Table 2). This suggests that these risks are primarily limited to state actors or sophisticated users with advanced knowledge in biology and machine learning, reducing both scale of risk and the “attack surface” that needs to be monitored (Rose et al., 2024; Ouagrham-Gormley, 2014; Dang et al., 2024).
@@ 4 Other potential threat models
Other, analyses highlight that LLMs may considerably or significantly uplift malicious users’ ability to develop harmful biological artifacts through improving laboratory experimentation and troubleshooting, and reducing barriers to biological work, for example by enabling technical coding for computational biology to be performed with non-technical, natural language (Rose et al., 2024). There are few empirical studies of this threat model, though one example is reported in OpenAI’s o1 system card which assesses model capabilities across a range of biological tasks, such as troubleshooting wet lab protocols and automating wet lab work. From these evaluations, OpenAI reports that "o1 can help experts with the operational planning of reproducing a known biological threat." However, considering other barriers to threat creation, they conclude that o1-preview and o1-mini pose “limited risk,” because the models do not provide biology experts with new knowledge, and "do not enable non-experts to create biological threats, because creating such a threat requires hands-on laboratory skills that the models cannot replace” (OpenAI, 2024c). With regards to AI tools assisting lab-based work, it is notable that existing risk frameworks focus on mitigating physical synthesis of threats, for example focusing on containment and control procedures that secure the handling and production of biological artifacts, rather than digital capabilities that support knowledge gathering, task automation, or planning (US Department of Health and Human Sciences, 2015).